OpenStudy (anonymous):

The perimeter of a rectangle is 66 m. If the width were doubled and the length were increased by 25 m, the perimeter would be 140 m. What are the length and width of the rectangle?

OpenStudy (anonymous):

2l+2w=66m 2l+25+4w=140m

OpenStudy (anonymous):

2w+2l=66 4w+2l+50=140 Thus 2w+2l=66 4w+2l=90 2w =24 w=12 l=21

OpenStudy (anonymous):

thank you!